OpenStudy (anonymous):

Find the derivative of f(x) = -cos (2x^2-3)

OpenStudy (anonymous):

Do you know how to use the chain rule?

OpenStudy (anonymous):

yeah yes i do

OpenStudy (anonymous):

Well your inner function is \(2x^2-3\) and your outer function is \(-\cos(x)\). There shouldn't be anything hard about this.

OpenStudy (anonymous):

do we have to use chain rule ? i'm confused on what we should do when with the -cos

OpenStudy (anonymous):

The - is just \(-1\cdot \cos(x)\). It's a constant. You can pull it out and put it back in after.

OpenStudy (anonymous):

oh really? but this have to use f(x+h)-f(x) first derivation ..

OpenStudy (anonymous):

You have to use this you mean? \[\Large \lim_{h\rightarrow 0}\frac{f(x+h)-f(x)}{h}\]

OpenStudy (anonymous):

You're not allowed to use the chain rule?

OpenStudy (anonymous):

Oh sorry, i just realised it only has to use chain rule. I make things go more complicated -_-

OpenStudy (anonymous):

So can you do it now?

OpenStudy (anonymous):

yeah the answer is sin(2x^2-3)(4x) right ?

OpenStudy (raden):

yeah, i agree that

OpenStudy (anonymous):

Oh thank you so much :D

OpenStudy (anonymous):

It's correct.